OpenStudy (anonymous):

K =1/3 (b – 2a) for b. Show work please.

OpenStudy (anonymous):

Multiply both sides by (b-2a) and let us see what you got?

OpenStudy (anonymous):

b - 2a is in the denominator or numerator??

OpenStudy (anonymous):

@kbreezy102 Is the equation: \[K = \frac{1}{3}*(b-2a)\]

OpenStudy (anonymous):

Or is it: \[K = \frac{1}{3*(b-2a)}\]

OpenStudy (anonymous):

its just 1/3(b-2a) the whole thing is not a fraction

OpenStudy (anonymous):

just the 1/3

OpenStudy (anonymous):

So the equation that wired has written is right??

OpenStudy (anonymous):

Can you draw the equation???

jimthompson5910 (jim_thompson5910):

It sounds like wired wrote the correct equation so you multiply both sides by 3 to start off to go from K = 1/3(b - 2a) to 3K = b - 2a Do you see what to do next?

OpenStudy (anonymous):

yes wired wrote the equation right

jimthompson5910 (jim_thompson5910):

So you just need to solve 3K = b - 2a for b

OpenStudy (anonymous):

subtract 2a?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

You can use the addition law of equality here..

jimthompson5910 (jim_thompson5910):

oops I thought you meant subtract -2a

jimthompson5910 (jim_thompson5910):

We have 3K = b - 2a add 2a to both sides to get 3K+2a = b - 2a+2a

OpenStudy (anonymous):

i mean add 2a?

jimthompson5910 (jim_thompson5910):

Yes that's better

OpenStudy (anonymous):

Yes now find for b..

jimthompson5910 (jim_thompson5910):

So adding 2a to both sides gives us 3K + 2a = b which becomes b = 3K + 2a

jimthompson5910 (jim_thompson5910):

The -2a+2a becomes 0a, which is 0 (and goes away)